OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. 3x – 3y = –8 x – 3y = –9 x – 3y = –8 3x + y = 9

OpenStudy (luigi0210):

You got this! You've done them before.

OpenStudy (anonymous):

i dont got this @Luigi0210

OpenStudy (luigi0210):

Yes you do! >.< Okay, first let's find the slope

OpenStudy (anonymous):

wait is it A??

OpenStudy (anonymous):

luigiiiiiiii

OpenStudy (luigi0210):

What? oh sorry about that. Hold on.

OpenStudy (luigi0210):

\[\frac{ 4-3 }{ 4-1 }=\]

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

i'm pretty sure the answer is a

OpenStudy (luigi0210):

Nope :/

OpenStudy (anonymous):

huh....

OpenStudy (anonymous):

please just tell me the answer luigi i've been working on this for like 3 hours straight ):

OpenStudy (luigi0210):

So we have the slope right? and points, so let's plug that into the point slope formula. \[y-3=\frac{ 1 }{ 3 }(x-1)\] Now distribute and solve: \[y-3=\frac{ 1 }{ 3 }x-\frac{ 1 }{ 3 }\] Add 3: \[y-3(+3)=\frac{ 1 }{ 3 }x-\frac{ 1 }{ 3 }+(\frac{ 9 }{ 3 })\] \[y=\frac{ 1 }{ 3 }x+\frac{ 8 }{ 3 }\] Now get it in standard form: ax+by=C \[-\frac{ 1 }{ 3 }x+y=\frac{ 8 }{ 3 }\] We have a negative fraction, and a cannot be negative nor a fraction so we have to get rid of it by multiplying it by -3 to both sides: \[-3(\frac{ -1 }{ 3 }x+y)=-3*(\frac{ 8 }{ 3 })\] That gives us: \[x-3y=-8\]

OpenStudy (anonymous):

awwww thanks luigi you're the bomb

OpenStudy (luigi0210):

^_^' Just here to help.

OpenStudy (anonymous):

you're awesome! :)